Question 11 Directions: Prepare the following strength solutions. Keep answer as a whole number. 1/4 strength Ensure 8 oz p.o. q4h for 24 hours How much Ensure will you use...express your answer in mLs. Question 12 Directions: Prepare the following strength solutions. Keep answer as a whole number. 14 strength Ensure 8 oz p.о. q4h for 24 hours How much water will you add...express your answer in mLs?

Show your work and answers for these questions
Answer #1

1/4 strength means= 8 ounces

Here one ounce= 30ml

So,

8ounce=8×30ml=240ml

Answer: 240ml

Question 12:

1/4 strength : 8 ounce

1ounce=30ml

So

8 ounce=8×30ml

=240ml

Answer=240ml.